# On Efficient Variants of Segment Anything Model: A Survey

### v1

@@ 4.3 Accuracy Comparison
Model $pt_{1}$ 1*$pt_{2}$ 3*$pt_{2}$ $pt_{1}$ 1*$pt_{2}$ 3*$pt_{2}$
SAM-H 53.6 55.6 67.5 60 59.8 65.0
SAMfast-H 52.9 54.5 67.4 53.6 57.2 60.6
SAM2-B+ 54.3 54.9 68.5 53.1 55.9 62.6
FastSAM 46.7 47.3 48.7 28.4 29.8 29.5
MobileSAM 48.6 49.2 59.1 45.6 47.2 50.0
EdgeSAM 47.2 46.8 61.6 45.2 46.5 54.4
EfficientSAM-Ti 43.8 44.6 61.6 25.4 26.9 52.5
EfficientSAM-S 41.1 39.4 60.1 52.3 53.1 66.3
RepViT-SAM 50.6 51.8 63.1 51.0 51.6 54.6
EfficientViT-SAM-L0 50.1 50.7 67.2 46.4 48.3 63.9
EfficientViT-SAM-XL1 54.3 55.7 70.2 53.5 54.4 65.7
SlimSAM-50 48.5 48.5 63.5 51.7 53.4 58.6
SlimSAM-77 47.2 47.2 61.1 48.6 50.4 55.3
TinySAM 42.4 42.1 65.1 50.0 51.0 60.9
NanoSAM 44.2 42.4 56.3 42.7 41.4 48.3
Table 5: The mIoU results on COCO and LVIS with points as prompts. P.S. $pt_{1}$ denotes the center point of the ground truth bounding box and $pt_{2}$ denotes points randomly sampled from ground truth masks.
COCO LVIS
Model $box_{1}$ $box_{2}$ $box_{1}$ $box_{2}$
SAM-H 77.4 77.8 78.0 79.1
SAMfast-H 77.3 77.7 77.0 78.0
SAM2-B+ 75.7 76.6 70.4 72.4
FastSAM 60.5 60.6 50.4 50.6
MobileSAM 73.4 74.0 74.4 74.8
EdgeSAM 75.9 76.2 74.4 74.8
EfficientSAM-Ti 72.6 73.3 72.9 74.5
EfficientSAM-S 69.1 69.0 74.7 75.5
RepViT-SAM 75.1 75.7 73.6 74.8
EfficientViT-SAM-L0 78.5 78.2 78.0 77.9
EfficientViT-SAM-XL1 79.9 79.8 79.9 79.7
SlimSAM-50 75.5 76.3 74.3 76.0
SlimSAM-77 74.4 75.2 72.7 74.5
TinySAM 73.7 73.9 73.6 74.1
NanoSAM 69.7 70.2 65.1 66.5
Table 6: The mIoU results on COCO and LVIS with boxes as prompts. P.S. $box_{1}$ denotes the ground truth bounding box and $pt_{2}$ denotes the tightest bounding box with respect to the ground truth mask.
To evaluate on the SegAny task, we adopt two types of points as prompts: 1) the center point of the ground truth bounding box, and 2) random points uniformly sampled from ground truth masks, following the setting in [131]. We evaluate the variants on COCO and LVIS, and the mean intersection over union (mIoU) is reported in Tab. 5. When prompted with the center point, SAM2-B+ and EfficientViT-SAM-XL1 achieve the highest mIoU of 54.3% on COCO, surpassing SAM-H with 53.6% mIoU, while SAMfast-H, also with 53.6%, exhibits the best performance among variants on LVIS. Under the setting of random point prompts, it is demonstrated that EfficientViT-SAM-XL1 performs better than SAM-H, particularly when prompted with 3 points, with an increase of 2.7% and 0.7%, respectively. From the perspective of datasets, we observe that the results on LVIS are generally lower than those on COCO, especially for FastSAM and EfficientSAM-Ti, whose accuracy decreases to below 30% on LVIS.
Moreover, we also evaluate the accuracy on SegAny task with two types of box prompts: 1) The ground truth bounding box and 2) The tightest bounding box corresponding to the ground truth mask, inspired by the experiments in [131, 159]. We reported the results of mIoU on COCO and LVIS which are illustrated in Tab. 6. We observe that EfficientViT-SAM-XL1 demonstrate the highest accuracy in every setting which exceeds SAM-H by 1.5%, 1.1%, 1.9% and 0.6%, respectively. SAMfast-H and EfficientViT-SAM-L0 also present performance close to SAM-H in the box-prompted segmentation tasks.
@@
For instance segmentation tasks, we adopt the ViTDet [59], YOLOv8 [53], GrounddingDINO [69], Detic [165] and H-Deformable-DETR [52] with Swin-L [71] as the object detectors which helps to generate bounding boxes of latent objects, with reference to [117, 164, 159]. We evaluate the average precision (AP) of all objects, as well as AP of small, medium, and large objects. The results are reported in Tab. 7, 8 and 9. Similar to the previous results, we find that on COCO dataset EfficientViT-SAM-XL1 always presents the highest AP with any of the detectors (except H-Deformable-DETR). Under the setting of equipping ViTDet as detector and testing on LVIS dataset, SAMfast-H surpass all other variants with AP of 44.5%.
@@ 5 Conclusion
In this survey, we have primarily discussed and evaluated the prominent works that focus on methods to efficiently segment anything and segment everything with reduced resource consumption and lower latency. For efficient SegAny tasks, most works adopt the approach of replacing either the image encoder or the entire architecture with a lightweight alternative, followed by training from scratch or through knowledge distillation. Other works aim to compress the original model by leveraging techniques such as quantization, pruning, or local optimization. For efficient SegEvery tasks, adopting an effective and efficient sampling strategy for prompt generation is essential. After reviewing these methods in detail, we have also outlined four potential future research directions that may drive new trends in this area. Additionally, we have evaluated the efficiency, accuracy, and the corresponding trade-offs of these models in a consistent environment, providing a fair and valuable comparison. Our analysis shows that some variants have already outperformed the original SAM in specific scenarios, and we believe their success will inspire further exploration and innovation in this field.

### v2

@@ 4.3 Accuracy Comparison
Table 5: Quantitative results of the accuracy of SegAny task (mIoU) on COCO and LVIS with points and boxes as prompts. For evaluation with points prompts, we select the center point of the ground truth bounding box ($pt_{1}$), and one or three randomly sampled points from ground truth masks (1*$pt_{2}$, 3*$pt_{2}$). For evaluation with boxes prompts, we adopt the ground truth bounding box ($box_{1}$), and the tightest bounding box calculated by the ground truth mask ($box_{2}$).
@@
Model $pt_{1}$ 1*$pt_{2}$ 3*$pt_{2}$ $box_{1}$ $box_{2}$ $pt_{1}$ 1*$pt_{2}$ 3*$pt_{2}$ $box_{1}$ $box_{2}$
SAM-H 53.6 55.6 67.5 77.4 77.8 60 59.8 65.0 78.0 79.1
SAMfast-H 52.9 54.5 67.4 77.3 77.7 53.6 57.2 60.6 77.0 78.1
SAM2-B+ 54.3 54.9 68.5 75.7 76.6 53.1 55.9 62.6 70.4 72.4
FastSAM 46.7 47.3 48.7 60.5 60.6 28.4 29.8 29.5 50.4 50.6
MobileSAM 48.6 49.2 59.1 73.4 73.9 45.6 47.2 50.0 74.4 74.8
EdgeSAM 47.2 46.8 61.6 75.9 76.2 45.2 46.5 54.4 74.4 74.8
EfficientSAM-Ti 43.8 44.6 61.6 72.6 73.3 25.4 26.9 52.5 72.9 74.5
EfficientSAM-S 41.1 39.4 60.1 69.1 69.0 52.3 53.1 66.3 74.7 75.5
RepViT-SAM 50.6 51.8 63.1 75.1 75.7 51.0 51.6 54.6 73.6 74.8
EfficientViT-SAM-L0 50.1 50.7 67.2 78.5 78.2 46.4 48.3 63.9 78.0 77.9
EfficientViT-SAM-XL1 54.3 55.7 70.2 79.9 79.8 53.5 54.4 65.7 79.9 79.7
SlimSAM-50 48.5 48.5 63.5 75.5 76.3 51.7 53.4 58.6 74.3 76.0
SlimSAM-77 47.2 47.2 61.1 74.4 75.2 48.6 50.4 55.3 72.7 74.5
TinySAM 42.4 42.1 65.1 73.7 73.9 50.0 51.0 60.9 73.6 74.1
NanoSAM 44.2 42.4 56.3 69.7 70.2 42.7 41.4 48.3 65.1 66.5
To evaluate on the SegAny task, we adopt two types of points as prompts: 1) the center point of the ground truth bounding box, and 2) random points uniformly sampled from ground truth masks, following the setting in [143]. We evaluate the variants on COCO and LVIS, and the mean intersection over union (mIoU) is reported in Tab. 5. When prompted with the center point, SAM2-B+ and EfficientViT-SAM-XL1 achieve the highest mIoU of 54.3% on COCO dataset, surpassing SAM-H with 53.6% mIoU, while on the LVIS dataset, SAMfast-H, with mIoU of 53.6%, exhibits the best performance among variants. Under the setting of random point prompts, it is demonstrated that EfficientViT-SAM-XL1 performs better than SAM-H on both datasets when prompted with 3 points, with the increase of 2.7% and 0.7%, respectively. From the perspective of datasets, we observe that the results on LVIS are generally lower than those on COCO.
Moreover, we also evaluate the accuracy on SegAny task with two types of box prompts: 1) The ground truth bounding box and 2) The tightest bounding box corresponding to the ground truth mask, instructed by the experiments in [143, 171]. We reported the results of mIoU on COCO and LVIS which are illustrated in Tab. 5. We observe that EfficientViT-SAM-XL1 demonstrate the highest accuracy in every box-prompted setting, which exceeds SAM-H by 2.5%, 2.0%, 1.9% and 0.6%, respectively. SAMfast-H and EfficientViT-SAM-L0 also present performance close to SAM-H in the segmentation tasks with box prompts.
@@
For instance segmentation tasks, we adopt the ViTDet [66], YOLOv8 [60], GrounddingDINO [79], Detic [177] and H-Deformable-DETR [59] with Swin-L [81] as object detectors to generate bounding boxes of latent objects, with reference to the experiments conducted in [129, 176, 171]. We evaluate the average precision (AP) of all objects, as well as AP of small, medium, and large objects. The results are reported in Tab. 6, 7 and 8. Similar to the previous results, we find that on COCO dataset EfficientViT-SAM-XL1 always presents the highest AP with any of the detectors (except H-Deformable-DETR). Under the setting of equipping ViTDet as detector and testing on LVIS dataset, SAMfast-H surpass all other variants with AP of 44.5%.
@@ 5 Conclusion
In this survey, we have primarily discussed and evaluated the prominent works that focus on methods to efficiently segment anything and segment everything with reduced resource consumption and lower latency. For efficient SegAny tasks, most works adopt the approach of replacing either the image encoder or the entire architecture with a lightweight alternative, followed by training from scratch or through knowledge distillation. Other works aim to compress the original model by leveraging techniques such as quantization, pruning, or local optimization. For efficient SegEvery tasks, adopting an effective and efficient sampling strategy for prompt generation is essential. After reviewing these methods in detail, we have also outlined five potential future research directions that may drive new trends in this area. Additionally, we have evaluated the efficiency, accuracy, and the corresponding trade-offs of these models in a consistent environment, providing a fair and valuable comparison. Our analysis shows that some variants have already outperformed the original SAM in specific scenarios, and we believe their success will inspire further exploration and innovation in this field.